OpenStudy (allieboo):

9 5/9 - 6 5/6 simplify if necessary im in fifth grade so i dont do negative numerals yet.

OpenStudy (amistre64):

fifthe grade? theres an age limit for this site by the way ....

OpenStudy (allieboo):

i dont care i need help

OpenStudy (amistre64):

in other words, its best if you dont post your age or grade.

OpenStudy (allieboo):

i didnt post my age

OpenStudy (amistre64):

fractions are just additions to whole numbers so we can rewrite this as 9 + 5/9 - 6 - 5/6 and try to work it from there

OpenStudy (amistre64):

5th grade is about 10 years old on average, just saying

OpenStudy (allieboo):

what if i got held back

OpenStudy (amistre64):

it still doesnt change the fact that its best not to publish that information

OpenStudy (allieboo):

i have to make the fractions denominators equa;

OpenStudy (allieboo):

equal

OpenStudy (amistre64):

9 + 5/9 - 6 - 5/6 9-6 + 5/9 - 5/6 3 + 5/9 - 5/6 you can, but i do it another way so that math solves it for me

OpenStudy (asnaseer):

\[9\frac{5}{9}-6\frac{5}{6}\]first turn both into improper fractions:\[\frac{86}{9}-\frac{41}{6}\]then find a common multiple of both denominators (9 and 6) which is 18 in this case and multiply numerator and denominator by the same number such that the denominator ends up with 18:\[\frac{2*86}{2*9}-\frac{3*41}{3*6}=\frac{172}{18}-\frac{123}{18}\]now you can combine the fractions as they have the same denominator:\[=\frac{172-123}{18}=\frac{49}{18}\]

OpenStudy (amistre64):

say its equal to some number (N) 3 + 5/9 - 5/6 = N then solve it easily by multiplication 6(3 + 5/9 - 5/6 = N) 18 + 5(2)/3 - 5 = 6N 3(18 + 10/3 - 5 = 6N) 54 + 10 - 15 = 18N 64 - 15 = 18N 54 - 5 = 18N 49 = 18N 49/18 = N is the number

OpenStudy (allieboo):

asnaseer i dont need the top part of your explination otherwise thnx